A furniture manufacturer can sell dining-room tables for $70 apiece. The manufacture‘s total cost consists of a fixed overhead of $8,000 plus production costs of $30 per table.


a. How many tables must be the manufacturer sell to break even?


b. How many tables must the manufacturer sell to make a profit of $6,000?


c. What will the manufacturer’s profit or loss if 150 tables are sold?



1
Expert's answer
2022-03-16T11:10:18-0400

Solution

Price"=70"

Let "T" represent the total number of tables sold.

Total revenue

"TR=70T"

Total cost "TC=Fixed~cost +Variable ~cost"

"TC=8,000+30T"


(a) To break even total revenue must equal total cost


"TR=TC"

"70T=8,000+30T"


solving for "T"

"T= 40" tables


(b) To make a profit of "6,000"


"TR=TC+6,000"

"70T=(8,000+30T)+6,000"


Solving for "T"

"T =350" tables


(c) Net gain from 150 tables

"TR=70T=70\times 150=10,500"

"TC=8,000+30\times150=12,500"


Net gain "=TR-TC"

"=10,500-12,500"

"=-2,500"


The manufacturer will have a loss of "2,500" if only 150 tables are sold.